Suppose that market for good X has a four-firm concentration ratio of 0.80. Having worked for the four largest firms in the industry, you know the sales for these four firms are given by $100,000, $125,000, $150,000, and $175,000. Based on this information, what are sales for the remaining firms in the industry?
Answer #1

Let the sales of remaining firms be x

Concentration ratio =

total sales of leading firms*100/total sales in the industry

Leading firms sales= (100000+125000+150000+175000)= 550000

=> 0.8 = 550000*100/(550000 + x)

440000 + 0.8 x = 55000000

x= 68200000

Sales of the remaining firms in the industry= 68200000